Required information

[The following information applies to the questions displayed below.] Kase, an individual, purchased some property in Potomac, Maryland, for $152,000 approximately 10 years ago. Kase is approached by a real estate agent representing a client who would like to exchange a parcel of land in North Carolina for Kase’s Maryland property. Kase agrees to the exchange. What is Kase’s realized gain or loss, recognized gain or loss, and basis in the North Carolina property in each of the following alternative scenarios? (Loss amounts should be indicated by a minus sign. Leave no answer blank. Enter zero if applicable.)

a. The transaction qualifies as a like-kind exchange and the fair market value of each property is $710,000.

      b. The transaction qualifies as a like-kind exchange and the fair market value of each property is $116,000.
Answer #1

a. (1) Amount realized from machine = $710,000

(2) Amount realized from boot (non like-kind property) = $0

(3) Total amount realized = (1) + (2) = $710,000

(4) Adjusted basis = $152,000

(5) Gain realized = (3) - (4) = $558,000

(6) Gain recognized = Lesser of (2) or (5) = $0

(7) Deferred gain = (5) - (6) = $558,000

Adjusted basis in new property = (1) - (7) = $152,000

Realized gain = $558,000

Recognised gain = $0

Adjusted basis in new property = $152,000

b. (1) Amount realized from machine = $116,000

(2) Amount realized from boot (non like-kind property) = $0

(3) Total amount realized = (1) + (2) = $116,000

(4) Adjusted basis = $152,000

(5) Loss realized = (3) - (4) = $(36,000)

(6) Loss recognized = Lesser of (2) or (5) = $0

(7) Deferred loss = (5) - (6) = $36,000

Adjusted basis in new property = (1) - (7) = $152,000

Realized loss = ($36,000)

Recognised loss = $0

Adjusted basis in new property = $152,000